OpenStudy (anonymous):

find zeroes of x^4+36x^2

OpenStudy (abb0t):

factor out x-squared and set equal to zero you should get x = 0 x = plus or minus 6.

OpenStudy (anonymous):

it cant be six because (x+6) (x-6) = x^2 -36 not +36 so i guess its only 0... thanks.

OpenStudy (abb0t):

but you're trying to find the zero's. Lol

hartnn (hartnn):

yeah, its not 6 x^2(x^2+36) = 0 x^2 = 0 , x^2 =- 36 x=0,0 , x=\(\sqrt{-6^2} \implies x=0,0,+6i,-6i \\where \:\: i=\sqrt{-1}\)